# Behavior of prediction performance metrics with rare events

### v1

@@ Abstract
Area under the receiving operator characteristic curve (AUC) is commonly reported alongside binary prediction models. However, there are concerns that AUC might be a misleading measure of prediction performance in the rare event setting. This setting is common since many events of clinical importance are rare events. We conducted a simulation study to determine when or whether AUC is unstable in the rare event setting. Specifically, we aimed to determine whether the bias and variance of AUC are driven by the number of events or the event rate. We also investigated the behavior of other commonly used measures of prediction performance, including positive predictive value, accuracy, sensitivity, and specificity. Our results indicate that poor AUC behavior—as measured by empirical bias, variability of cross-validated AUC estimates, and empirical coverage of confidence intervals—is driven by the minimum class size, not event rate. Performance of sensitivity is driven by the number of events, while that of specificity is driven by the number of non-events. Other measures, including positive predictive value and accuracy, depend on the event rate even in large samples. AUC is reliable in the rare event setting provided that the total number of events is moderately large.
Funding
The authors disclosed receipt of the following financial support for the research, authorship, and/or publication of this article: Research reported in this publication was supported by the National Institute of Mental Health of the National Institutes of Health under award numbers R01 MH125821, U19-MH099201, and U19-MH121738.
Acknowledgments
The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.
Declaration of competing interest
The authors declared no potential conflicts of interest with respect to the research, authorship, and/or publication of this article.
Ethical approval and informed consent
No new data were created for this study. The data analyzed in this study were obtained under approval from the appropriate institutional review boards for the participating health systems.
Data availability
The datasets generated and analyzed during this study are not publicly available because they contain detailed information from the electronic health records in the health systems participating in this study and are governed by the Health Insurance Portability and Accountability Act (HIPAA). Data are however available from the authors upon reasonable request, with permission of all health systems involved and fully executed data use agreement.
@@ 1 Introduction
Clinical prediction models can be used to guide clinical decision-making through early warning systems (Patton and Liu, 2023), identify patients at risk for an adverse outcome (Simon et al., 2018), obtain accurate prognoses and diagnoses (Obermeyer and Emanuel, 2016), and better allocate health resources (Chen and Asch, 2017). When the outcome of interest is binary, prediction performance can be measured in a variety of ways, including sensitivity and specificity (see, e.g., Metz, 1978), Brier score (see, e.g., Steyerberg et al., 2001), and the area under the receiver operating characteristic curve (AUC; see, e.g., Hanley and McNeil, 1982). Often, assessing several prediction performance metrics can provide a full picture of model performance, but focus on any one metric depends on how the model will ultimately be used (Cook, 2007; Pepe et al., 2007; Janket et al., 2007).
Estimating and reporting the performance of risk prediction models in the rare event setting presents several challenges. First, even with a large number of observations, the absolute number of events may be small. In predictive modeling of binary outcomes, the number of events, not just the total sample size, is of importance (Vergouwe et al., 2005). Additionally, a small event rate impacts the interpretation of some common measures of predictive model performance because the overall event rate impacts the behaviors of these measures (Lever et al., 2016; Saito and Rehmsmeier, 2015). There has been some recent concern that AUC, and receiver operating characteristic curves more generally, are misleading or uninformative in the rare event setting (Saito and Rehmsmeier, 2015; Lever et al., 2016; Steyerberg et al., 2018; Adhikari et al., 2021; Bokhari, 2023). In this manuscript, we aimed to determine whether a low event rate regardless of sample size or a small number of events alone leads to either instability in AUC estimates or AUC estimates that do not reflect the true AUC.
Our motivation for considering AUC in the rare-event setting is that the AUC, among other prediction performance metrics, is often used in the literature to describe the performance of suicide risk prediction models. In our collaborations developing suicide risk prediction models (Simon et al., 2018; Coley et al., 2021; Shortreed et al., 2023), the per-visit 90-day suicide attempt rate is close to 0.65% for visits to mental health specialty clinics and 0.33% for general medical visits with mental health diagnoses; suicide deaths are even more rare, approximately 0.02% and 0.01% for mental health specialty visits and general medical visits, respectively. Despite this rarity, suicide was the 10th leading cause of death in the United States in 2019; there were over suicide deaths, representing suicide deaths per population (Xu et al., 2021). Predicting suicide deaths and attempts in order to identify individuals at risk—so that these individuals can be offered interventions—is clearly of clinical importance. A 2019 systematic review found a total of 64 unique suicide prediction models (Belsher et al., 2019). While suicide attempt is rare, the total sample size for developing these prediction models can be large. Thus, having a clear understanding of the behavior of AUC and other prediction metrics in the rare-event setting is important.
While the focus of the experiments presented here is AUC, we also investigated the behavior of multiple other metrics: sensitivity, specificity, positive predictive value (PPV) and negative predictive value (NPV), the F score (Van Rijsbergen, 1974), classification accuracy, and the Brier score. Each of these metrics depends differently on event rate, the number of events, and the number of non-events, and thus, may behave differently than AUC.
The remainder of the article is organized as follows. In the next section, we provide explicit definitions of the prediction performance metrics we consider. In Section 3, we describe our simulation methods, and present the results in Section 4. We provide concluding remarks in Section 5. Results of additional simulations, and results for several performance metrics, can be found in the Supplementary Material.
2 Prediction performance metrics: notation and definitions
@@
We suppose that we have independent and identically distributed observations drawn from a distribution , where the data unit is ; is a covariate vector and is the binary outcome of interest. Suppose further that we have a prediction model that has been trained on an independent sample drawn from (e.g., a training sample or a set of folds in a cross-validation procedure) and that returns a predicted probability of the outcome based on the observed covariates . Then, the AUC of prediction model with respect to distribution is
where and are independent draws from . In other words, the AUC is the probability that the predicted probability is correctly ordered lower for a randomly sampled non-event than for a randomly sampled event. The AUC can take values between 0.5 and 1; an AUC of 0.5 implies that the prediction model performs no better than assigning predicted probabilities at random (Bradley, 1997), while an AUC of one implies that the predicted probability of an event is always larger than the predicted probability of a non-event.
Next, we define sensitivity, specificity, PPV and NPV, the F score (Van Rijsbergen, 1974), classification accuracy, and the Brier score. All metrics besides the Brier score are defined relative to a threshold, , of predicted risk. The sensitivity (sens) and specificity (spec) of given threshold with respect to are
that is, sensitivity is the probability that the model-predicted probability of an event is higher than threshold given that the observation truly is an event and specificity is the probability that the model-predicted probability of an event is lower than threshold given that the observation truly is a non-event.
The positive and negative predictive values of given threshold relative to are
in other words, PPV is the probability that an observation is an event given the model-predicted probability of an event is greater than threshold and NPV is the probability that an observation is a non-event given the model-predicted probability of an event is less than threshold .
The F score of model given threshold with respect to trades off sensitivity and PPV; where the amount of the trade off is dictated by a constant, :
Two common choices are and , resulting in
The score weights sensitivity and PPV equally, while the score places higher weight on PPV; this can be useful in settings where there is a high cost to intervening based on . In many rare outcome settings the potential interventions under consideration are relatively high cost, thus putting a larger weight on PPV assures that we will be intervening on individuals at higher risk, potentially at a cost of missing some events.
The classification accuracy of given threshold with respect to is
where is the indicator function, i.e., if is true and otherwise. In other words, accuracy is the probability that the model-predicted event status defined using threshold is equal to the true event status, giving equal weight to events and non-events (Jiang et al., 2021).
Finally, the Brier score of with respect to is
where is shorthand for the expectation under . In other words, the Brier score is the mean squared error of .
3 Simulation methods
@@
We now outline the methods used to investigate the behavior of each these performance metrics in rare-event settings, with both large and small sample size.
3.1 Data
@@
There has been growing interest in estimating risk prediction models for self-harm and suicide, important and rare health outcomes. This work evaluating the behavior of prediction model performance metrics is motivated by our collaborations estimating suicide risk prediction models using health records data (see, e.g., Simon et al., 2018; Coley et al., 2021). To maintain both the complex outcome-predictor relationships and the complex correlation structure seen between predictors in our motivating dataset, we used a plasmode simulation approach (see, e.g., Schreck et al., 2024). Specifically, we constructed a “parent” dataset with independent observations by randomly sampling one mental health visit per patient from 25 million outpatient mental health visits across seven Mental Health Research Network sites (https://mhresearchnetwork.org). These visits were made by patients aged 13 years or older and occurred between January 1, 2009 and September 30, 2017 (Coley et al., 2021). For patients with an event at any visit—defined as non-fatal or fatal self-inflicted injury or poisoning within 90 days of a mental health visit—a visit with an event was sampled. The event rate in the resulting dataset, , was 0.92%. There were 149 predictors, consisting of demographic characteristics, mental health and substance abuse diagnoses (including past self-inflicted injuries or poisonings), other medical diagnoses, past inpatient or emergency mental health care, dispensed mental health medications, the patient health questionnaire (PHQ) 8-item score (PHQ-8) representing depressive symptoms, and the PHQ 9th item response indicating thoughts of death or self-harm (Coley et al., 2021; Simon et al., 2018). All predictors except patient age in years and PHQ-8 score (0–24) were coded as binary indicators; the PHQ item 9, which has 4 possible responses (0,1,2,3), was coded using 3 categorical variables indicating each possible non-zero response.
3.2 Simulation scenarios
@@
In our simulations, we bootstrap sampled the covariates with replacement from the parent dataset with observations. Then we generated outcomes according to a fitted regression model estimated in that original dataset (see, e.g., Franklin et al., 2017) with varying event rates. To do this, we sampled three datasets of size = 1 million with event rates %, %, and %. In each of these data sets, we fit logistic regression, ridge regression (Hoerl and Kennard, 1970), and random forest models (Breiman, 2001). These became three outcome-generating functions for our simulations for the three event rates. The tuning parameters used to fit each outcome model are provided in Table S1; 10-fold cross-validated estimates of the performance metric values on the appropriate dataset are presented in Table 1.
@@
To investigate the behavior of prediction performance metrics, we varied both the event rate (keeping sample size fixed) and the absolute number of events (increasing sample size appropriately as the event rate decreases). We varied and considered the three event rates defined above. For each scenario, we first bootstrap sampled with replacement covariates for observations (for a training dataset) and, separately, covariates for 1.6 million observations (for an evaluation dataset). The evaluation dataset is a large sample of completely independent data; performance in this dataset is intended to capture the true model performance in the population. Once we sampled covariates for each of the datasets (training, evaluation), using each of the three algorithms we generated outcomes from the appropriate model to obtain each event rate. For each scenario defined by event rate, sample size, and algorithm, we generated 2500 random datasets following the procedure described above (i.e., we conducted 2500 simulation iterations).
3.3 Developing clinical prediction models
@@
Estimation of prediction models within each training set used the same three prediction algorithms described above for generating outcomes: logistic regression, ridge logistic regression, and random forests. While logistic regression is a common prediction approach in binary outcome settings, in some rare-event cases, it may be unstable, particularly with a large number of predictors. Ridge regression was selected as an example of a penalized regression method that might be used in a setting with a large number of predictors (Hoerl and Kennard, 1970). Random forests were selected as an example of a more complex prediction modeling approach that allow for interactions between predictors without being explicitly coded, in contrast to logistic or ridge regression. We used probability trees (Malley et al., 2012) within random forests. In our simulations, we did not include any interactions between our predictors for logistic regression or ridge regression.
Ridge regression and random forest require the specification of several key tuning parameters: the regularization parameter for ridge, and the minimum node size, number of predictors, , considered at each split, the depth of tree, and number of trees for random forests. For random forests, we fixed (the square root of the number of predictors) and allowed tree depth to be controlled by minimum node size. We selected global tuning parameters that were used in all simulation replicates in a small simulation study with 100 replicates. In this small simulation study, for each simulation replication, we used 10-fold cross-validation to obtain optimal tuning parameters. Details on this simulation are provided in Section S1.5 of the Supplementary Materials. The final tuning parameters used for each algorithm in the main simulations presented here are provided in Table S2. We used global tuning parameters, because the goal of our simulation study was not to optimize the prediction performance of a given model but rather to understand how the behavior of prediction model performance measures depended on event rate and absolute number of events. We did not need the tuning parameters to be the optimal values for each sampled replication-specific training set, since we were interested in the performance of a model estimated with particular values of tuning parameters.
3.4 Assessing clinical prediction model performance
@@
To provide an accurate estimate of prediction performance based solely on the training data, we used cross-validation. The training set was divided into ten folds stratified by the outcome, and 10-fold cross-validation was performed to estimate AUC, sensitivity, specificity, PPV, NPV, accuracy, , , and Brier score. At the smaller training set sizes and event rates, on occasion folds had no events despite using outcome-stratified sampling (i.e., there were fewer than 10 events in the sampled training data). These folds without events were excluded during cross-validation, since at least one true event was required to obtain an estimate of the hold-out fold AUC, and AUC was the primary prediction performance metric of interest. The metrics requiring thresholds (all except AUC and Brier score) were calculated at each of the , , and percentiles of predicted probabilities from the training folds.
@@
We used empirical standard errors (ESEs) to describe the variability of performance measure metrics. Code implementing asymptotic standard errors (ASEs) were not available for many metrics. While the ESE is not available in a data analysis, in the simulation setting it provides a best-case benchmark for variance estimation if we were able to consistently estimate the standard error for each performance metric. For each performance metrics the ESE was defined as the standard deviation of the point estimates across the 2500 simulation replications. For AUC, in addition to the ESE, we calculated influence curve based confidence intervals (asymptotic standard errors, or ASEs) (LeDell et al., 2015), as code was readily available.
@@
We used the evaluation dataset to estimate the true prediction performance of each of the three algorithms in our population of interest. We fit each of the three algorithms on the entire training set in each simulation replication to obtain a fitted model and corresponding thresholds of predicted probabilities , , and . We computed all prediction performance metrics on the evaluation dataset using this fitted model and thresholds for each replication. The final “true" value of each prediction performance metric was then defined as the average of the evaluation-set values across the 2500 simulation replications.
@@
For each performance metric and each simulation scenario, we measured reliability by calculating the empirical bias, variance, and mean squared error (MSE) of the estimated performance metric, using the evaluation-estimated mean performance metric as the gold standard. We calculated empirical coverage of the 95% confidence intervals based on the ASE (for AUC) and on the ESE (for all metrics), comparing the confidence interval from each replication with the evaluation-estimated mean performance metric. The computation of these reliability measures is provided in Table 2.
@@
3.5 Computing
All simulations and analyses were conducted in R version 4.1.1 (R Core Team, 2013). Logistic regression was implemented using the base R stats package, penalized regression using the package glmnet (version 4.1-4) (Friedman et al., 2010), and random forests using the package ranger (version 0.14.1) (Wright and Ziegler, 2017). Influence curve-based confidence intervals for cross-validated AUC were obtained using the cvAUC package (LeDell et al., 2022). The foreach package was used to implement parallel computing (Daniel et al., 2022). For reproducibility, each replication of the simulations used a unique random number generation seed from a pre-defined sequence of random seeds.
4 Simulation results
We define the effective sample size for a prediction metric as the amount of information for estimating that metric in the entire training sample. For sensitivity, the effective sample size is the number of events in the training sample; for specificity, it is the number of non-events in the training sample. For AUC, the effective sample size is the minimum of the number of events and non-events; because the event rate in our scenario is less than 50%, the effective sample size is the number of events.
The true value of each performance measure depends on the sample size, event rate, and model-estimation algorithm. In Table 3, we show the true values of AUC in all settings; we present the true values for other estimands in Table LABEL:tab:all_true_values_test.

### v2

@@ Abstract
Objective Area under the receiving operator characteristic curve (AUC) is commonly reported alongside prediction models for binary outcomes. Recent articles have raised concerns that AUC might be a misleading measure of prediction performance in the rare event setting. This setting is common since many events of clinical importance are rare. We aimed to determine whether the bias and variance of AUC are driven by the number of events or the event rate. We also investigated the behavior of other commonly used measures of prediction performance, including positive predictive value, accuracy, sensitivity, and specificity.
Study Design and Setting We conducted a simulation study to determine when or whether AUC is unstable in the rare event setting by varying the size of datasets used to train and evaluate prediction models. This plasmode simulation study was based on data from the Mental Health Research Network; the data contained 149 predictors and the outcome of interest, suicide attempt, which had event rate 0.92% in the original dataset.
Results Our results indicate that poor AUC behavior—as measured by empirical bias, variability of cross-validated AUC estimates, and empirical coverage of confidence intervals—is driven by the number of events in a rare-event setting, not event rate. Performance of sensitivity is driven by the number of events, while that of specificity is driven by the number of non-events. Other measures, including positive predictive value and accuracy, depend on the event rate even in large samples.
Conclusion AUC is reliable in the rare event setting provided that the total number of events is moderately large; in our simulations, we observed near zero bias with 1000 events.
Plain Language Summary Predicting self-harm or suicidal behavior is medically important for guiding clinicians in providing care to patients. Several research teams have developed and evaluated suicide risk prediction models based on health records data. Part of evaluating these models is calculating area under the receiver operating characteristic curve (AUC) and other prediction performance metrics. Self-harm and suicide are rare events. Recent research has raised concerns with using AUC in rare-event settings. We aimed to determine whether having a sufficiently large dataset could remove these concerns. In our experiments, we found that AUC can be used without concern in settings with 1000 events or more. Thus, AUC is a valid measure of suicide risk prediction model performance in many large healthcare databases.
Keywords: Prediction, classification, machine learning, area under the receiver operating characteristic curve, model evaluation, rare outcome.
What is new
Key findings
AUC and other binary-outcome prediction performance metrics can be used reliably in rare-outcome settings so long as the minimum class size is large enough.
Multiple prediction performance metrics should be reported in binary outcome settings, along with estimates of variability.
What this adds to what is known?
We provide evidence that the bias and variance of AUC, sensitivity, and specificity are driven by effective sample size (number of events for sensitivity; number of non-events for specificity; and the minimum of these two for AUC). In our simulations, we found that 1000 events was sufficient to have near zero bias for AUC and near nominal confidence interval coverage.
We provide evidence that the bias and variance of other binary outcome prediction performance metrics also depends on the event rate.
What is the implication and what should change now?
Our results imply that in settings with rare events but large sample sizes—for example, predicting suicide attempt using millions of health care visits—AUC and other prediction performance metrics should reliably capture prediction performance. These measures should be used when scientifically relevant.
@@ 1 Introduction
Clinical decision-making can be guided by prediction models that identify people at risk for negative health outcomes [1, 2, 3, 4]. Assessing model performance across many metrics is important. Selecting any one particular metric depends on how the model will be used in care [5, 6, 7]. Sensitivity and specificity [8], Brier score [9], and the area under the receiver operating characteristic curve [10, AUC;] are all useful performance metrics when considering binary outcomes. Estimating and reporting prediction model performance in the rare event setting presents several challenges. First, even with a large number of observations, the absolute number of events may be small. In training predictive models, the number of events, not just the total sample size, is of importance [11]. Second, a small event rate impacts the behavior of some performance measures [12, 13].
AUC is often reported in the scientific literature to describe the performance of risk prediction models, as we and others have done when estimating models to predict suicide attempt [14, 15, 4, 16, 17, 18, 19, 20, 21, 22, 23]. In our prior work, the per-visit 90-day suicide attempt rate was 0.62% for visits to mental health specialty clinics and 0.26% for general medical visits with mental health diagnoses; suicide deaths were even more rare, approximately 0.02% and 0.01%, respectively [4]. Recently, several studies have argued that AUC is a misleading and inappropriate measure of prediction model performance in the rare event setting [13, 12, 24, 25, 26]. In practice, when using large sample sizes, even with very rare events, AUC estimates of suicide risk prediction model performance have aligned with other performance metrics; in other words, multiple metrics yield similar conclusions about model adequacy and strength of risk discrimination [14, 18, 23]. This is contrary to concerns raised in recent literature [12, 25]. It is essential for researchers and decision makers to have a clear understanding of the statistical properties of performance measures in rare-outcome settings, including bias (including the potential for optimism) and stability (i.e., estimator variability), as many prediction models are developed to predict rare outcomes. In this paper we examine bias and variance of AUC and other performance metrics in a variety of settings, varying both the sample size and the event rate.
2 Materials and Methods
2.1 Prediction performance metrics: notation and definitions
@@
We assume independent and identically distributed observations drawn from a distribution , with data unit , covariate vector and binary outcome . We further assume a prediction model, , trained on a sample from , such that returns a predicted probability of given . The AUC of is
where and are independent draws from . In other words, AUC is the probability that a randomly sampled non-event has a lower predicted probability than a randomly sampled event. We define other common performance metrics in Table 1.
| Metric | Definition |
| Sensitivity | |
| Specificity | |
| Positive predictive value | |
| Negative predictive value | |
| F0.5 score | |
| F1 score | |
| Classification accuracy | |
| Brier score |
2.2 Simulation methods
@@
We examined the behavior of AUC and the performance metrics in Table 1 in a variety of rare-event settings, with both large and small sample size.
2.2.1 Data
@@
To maintain both the complex outcome-predictor relationships and the complex correlation structure among predictors in real-world data [4, 20], we used a plasmode simulation approach [27]. We obtained a dataset with independent observations by randomly sampling one mental health visit per patient from across seven Mental Health Research Network sites (https://mhresearchnetwork.org). These visits were made by patients aged 13 years or older and occurred between January 1, 2009 and September 30, 2017 [20]. For patients with a suicide attempt (fatal or non-fatal) during this time, we sampled a mental health visit within 90 days; for those without a suicide attempt, a randomly sampled mental health visit was selected. The event rate in , , was 0.92%. We considered 149 predictors, including demographic characteristics, mental health and substance abuse diagnoses (including past self-inflicted injuries or poisonings), other medical diagnoses, past inpatient or emergency mental health care, dispensed mental health medications, the patient health questionnaire (PHQ) 8-item score (PHQ-8) representing depressive symptoms, and the PHQ 9th item response indicating thoughts of death or self-harm [4]. All predictors except patient age in years and PHQ-8 score (0–24) were coded as binary indicators. The PHQ item 9, which has 4 possible responses (0,1,2,3), was coded using 3 binary variables indicating each possible non-zero response.
2.2.2 Simulation scenarios
@@
We used three types of outcome-generating functions for our simulations: logistic regression, ridge regression [28], and random forest models [29]. Outcome-generating models for event rates of 0.92%, 0.46%, and 1.84% were created by sampling 1 million visits from with and without events to match the desired event rate. The tuning parameters used for each outcome-generating model are provided in Table S1; 10-fold cross-validated estimates of performance of these models is presented in Table 2.
@@
To investigate the behavior of prediction performance metrics, we varied both (i) the event rate, keeping sample size fixed, and (ii) the absolute number of events, increasing sample size appropriately as the event rate decreases. We varied and considered the three event rates defined above. For each simulation iteration we generated a training data set by sampling covariates for observations using bootstrap sampling with replacement from , and generated outcomes according to each prediction model [30, see, e.g.,]. We then repeated this procedure once for each sample size, event rate, and outcome-generating model to generate an evaluation data set of 1.6 million observations. Performance in this evaluation dataset, a large independent sample, is intended to capture true model performance. For each scenario, defined by event rate, sample size, and algorithm, we conducted 2500 simulation iterations.
2.2.3 Developing clinical prediction models
@@
For each simulation iteration we estimated logistic regression, ridge logistic regression, and random forests prediction models. While logistic regression is a common prediction approach in binary outcome settings, in some rare-event cases, it may be unstable, particularly with a large number of predictors. Ridge regression was used as a penalized logistic regression method [28]. Random forests with probability trees [31] was selected as an example of a more complex prediction modeling approach that estimated interactions between predictors automatically from the data. We did not include any interactions between predictors in models estimated with logistic regression or ridge regression. We used global tuning parameters in model estimation for each simulation iteration because our goal was to understand how the behavior of prediction model performance measures depended on event rate and absolute number of events, not estimate the optimal model for each data set. For information on tuning parameters used for the simulations, see Supplementary Materials Section S1.5.
2.2.4 Assessing clinical prediction model performance
@@
We used 10-fold outcome-stratified cross-validation to estimate performance metrics for each model on the training dataset. At smaller training set sizes and event rates, some folds had no events despite using outcome-stratified sampling (i.e., there were fewer than 10 events in the sampled training data). Folds without events were excluded from calculation of prediction performance. We used the , , and percentiles of predicted probabilities from the training fold for metrics requiring thresholds (all except AUC and Brier score).
@@
Code implementing asymptotic standard errors (ASEs) was not available for many metrics; thus, we used empirical standard errors (ESEs) to describe variability. While the ESE is not available in a data analysis, in simulations it provides a best-case benchmark for variance estimation. The ESE was defined as the standard deviation of estimates across the 2500 simulation replications. For AUC, in addition to the ESE, we calculated ASEs, as code was readily available [32, 33].
@@
We used the evaluation dataset to describe the true prediction performance in our population, computing all performance metrics using the fitted model and thresholds from each replication. The final true population value of each performance metric was defined as the average of the evaluation-set values across the 2500 simulation replications.
@@
We measured reliability for the performance metrics by calculating the empirical bias, variance, and mean squared error (MSE) of the estimated performance metric across the training-set estimated metrics. The evaluation-estimated mean performance metric was treated as the gold standard, i.e., the true performance (see Table 3). We calculated empirical coverage of 95% confidence intervals (CIs) using the ASE for AUC and the ESE for all metrics, comparing the CI from each replication with the evaluation-estimated mean performance metric.
@@
2.2.5 Computing
All simulations were conducted in R version 4.1.1 [34]. Logistic regression was implemented using the base R stats package, ridge regression using glmnet (version 4.1-4) [35], and random forests using ranger (version 0.14.1) [36]. ASE-based CIs for cross-validated AUC were obtained using cvAUC [33]. The foreach package was used to implement parallel computing [37].
3 Results
We define the effective sample size (ESS) as the amount of information in the entire training sample used to estimate a metric. Therefore, for sensitivity, the ESS is the number of events in the training sample; for specificity, it is the number of non-events. For AUC, the ESS is the minimum of the number of events and non-events; because we consider event rates less than 50%, in our case the ESS is the number of events. The true value of each performance measure depends on the sample size used to estimate the model, event rate, and model-estimation algorithm. The true values for AUC are shown in Table 4; for other estimands see Table LABEL:tab:all_true_values_test.